OpenStudy (anonymous):

What is the area of a square with side lengths of 7 feet? A. 28 ft B. 28 ft2 C. 49 ft2 D. 49 ft

OpenStudy (anonymous):

Multiply 7 * 7

OpenStudy (anonymous):

No, it's 7^2

OpenStudy (anonymous):

\(A = s^2\) \(A = 7^2\) \(A = 7 \times 7\)

OpenStudy (anonymous):

Oops, I meant 7^2.

OpenStudy (anonymous):

it equals 49

OpenStudy (anonymous):

Yes, that's our answer.

OpenStudy (anonymous):

thank u guys so much